OpenStudy (anonymous):

Given log base a 3=x and log base a 2=y, Write log base a 54/a^2 in terms of x and y

OpenStudy (anonymous):

\[\log_{a}3=x, \log_{a}2=y, Write \log_{a}\frac{ 54 }{ a^2 } \] in terms of x and y

OpenStudy (anonymous):

Medal and fan

OpenStudy (welshfella):

Hint Factor 54 54 = 2*3^3

OpenStudy (welshfella):

Also log a a^2 = 2